Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

FAN AND MEDAL A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate cookie, 3 almond cookies, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking 1 almond cookie followed by another almond cookie, and show the equation used.

OpenStudy (anonymous):

I think it is 3/6 * 2/5 = 6/30 =1/5

OpenStudy (anonymous):

I got the same thing so I think it is correct see.. 3/6 * 2/5 = 6/30 =1/5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!